Instructions: Solve each system by the elimination method.
What is the solution for 5x-y=5
-5x+2y=0

Answer by checkley71(8405)

5X-Y=5
-5X+2Y=0 ELIMINATE X BY ADDITION
---------------------------------------------
Y=5
THEN
5X-5=5
5X=10
X=2 SOLUTION
PROOF
5*2-Y=5
-Y=5-10
-Y=-5
Y=5 SOLUTION